OpenStudy (anonymous):

solve the equation 3x-7=5x+19

OpenStudy (anonymous):

x=-13

OpenStudy (alexwee123):

3x-7=5x+19 -7=2x+19 -7-19=2x -26=2x x=-13

OpenStudy (anonymous):

Subtract 3x both the sides: \[3x - 7 - 3x = 5x + 19 - 3x\] \[-7 = 2x + 19\] Subtract 19 both the sides now: \[-7 - 19 = 2 x + 19 - 19\] \[-26 = 2x\] Now divide by 2 both the sides to find x: \[\frac{2x}{2} = \frac{-26}{2} \implies \color{green}{x = -13}\]

OpenStudy (anonymous):

thank you